OpenStudy (anonymous):

If I run about 10 miles per hour and every 20 mintues I lose .10 miles how many miles do I have in 4 hours

OpenStudy (anonymous):

btw Its POINT ten miles and I know I spelled minutes wrong .

OpenStudy (anonymous):

If in 20 minutes, you lose 0.10 miles. In one hour, you lose, 60mins/20mins * 0.10 = 0.30 miles. So in one hour, you run 10 miles but lose 0.30, in total you run only 10-0.30 = 9.7miles In 4 hours, you run 9.7*4= 38.8 miles

OpenStudy (anonymous):

what about .2

OpenStudy (anonymous):

what .2? i don't see any .2 at all.

OpenStudy (anonymous):

what if you changed .10 to .2

OpenStudy (anonymous):

same method. If you lose .2 miles in 20 minutes. In one hour, u lose .2 *3 = .6 miles. In one hour, instead of running 10 miles, you actually run 9.4 miles. In 4 hours, 9.4 *4 = 37.6 miles

OpenStudy (anonymous):

how bout + .10

OpenStudy (anonymous):

same method. In 20 minutes you gain 0.1 miles. In one hour u gain 0.3 miles In one hour total, 10+0.3= 10.3miles in 4 hours, 10.3*4 = 41.2 miles

OpenStudy (anonymous):

how bout x .5